# Offline Clustering of Linear Bandits: Unlocking the Power of Clusters in Data-Limited Environments

### v1

@@ Appendix B Technical Lemmas
Lemma B.2.
Let $\bm{a}_{s},n\geq 1$ be generated sequentially from a random distribution such that $||\bm{a}||_{2}\leq 1$ and $\mathbb{E}[\bm{a}\bm{a}^{\top}]$ is full rank with minimal eigenvalue $\lambda_{a}>0$. Let $M_{n}=\sum_{s=1}^{n}\bm{a}_{s}\bm{a}_{s}^{\top}$. Then event
@@
$\lambda_{\min}(M_{n})\geq\frac{1}{2}\lambda_{a}n,\ \forall n\geq\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8d}{\lambda_{a}^{2}\delta}\right)$
@@
When $n\geq\frac{512d}{\gamma^{2}\lambda_{a}}\log(\frac{2U}{\delta})$ and $\lambda\leq d\log(1+\frac{n}{\lambda d})+2\log\left(\frac{2U}{\delta}\right)$,
$\frac{\sqrt{d\log(1+\frac{n}{\lambda d})+2\log(\frac{2U}{\delta})}+\sqrt{\lambda}}{\sqrt{\lambda_{a}\lambda+\lambda_{a}n/2}}\leq\frac{\gamma}{4}.$
@@
The proofs of the above lemmas can be found in Li and Zhang [2018] and Wang et al. [2025]. Specifically, Lemma B.1 corresponds to Lemma 1, Lemma B.3 to Lemma 9, and Lemma B.4 to Lemma 10 in Li and Zhang [2018]. Lemma B.2 corresponds to Lemma 7 in Li and Zhang [2018] and Lemma B.2 in Wang et al. [2025].
Appendix C Additional Experiments
C.1 Basic Setup and Dataset Description
[…]
@@ C.1 Basic Setup and Dataset Description
Dataset. (1). Synthetic Dataset: The user preference vectors $\bm{\theta}_{u}$ and item feature vectors $\bm{a}_{u}^{i}$ are in $d=20$-dimensional space, with each entry drawn from a standard Gaussian distribution and normalized to satisfy $\|\bm{\theta}_{u}\|_{2}=1$. (2). Real-world Datasets: We use two real-world recommendation datasets: We extract $1k$ users and $1k$ items with the highest interaction frequencies, then apply Singular Value Decomposition (SVD) to extract user preference vectors. For each user and each $i$, the action set $\mathcal{A}_{i}$ consists of 20 items, where each item’s feature vector is independently sampled from a sub-Gaussian distribution and then normalized to unit norm, following Li et al. [2019].
C.2 Benchmark Methods and Their Descriptions
To ensure fair comparisons, we adapt several online learning algorithms to the offline setting, modifying them as needed to work with offline datasets. Additionally, we propose an enhanced version of DBSCAN Schubert et al. [2017], tailored specifically to our environment. Compared to KMeans, DBSCAN is more suitable for our framework due to its parameterization (neighborhood radius and minimum samples), which aligns better with our requirements. For learning methods, We evaluate the following benchmarks:
  • SCLUB: A method that employs a hierarchical clustering structure to model users and enhances learning efficiency by optimizing cluster centers.
  • LinUCB: A linear regression-based Upper Confidence Bound (UCB) approach that dynamically balances exploration and exploitation for model selection.
@@ C.2 Benchmark Methods and Their Descriptions
  • LinUCB_Ind: An independent variant of LinUCB that operates without leveraging any clustering information. This method applies the LinUCB approach independently to each user.
@@
  • CLUB: A cluster-based multi-armed bandit method that groups users from a complete graph according to their characteristics to improve model selection.
  • DBSCAN_Improve: Similar to CLUB, but uses DBSCAN for the clustering stage.
  • XMeans_Improve: Similar to CLUB, but uses XMeans for the clustering stage.
Figure 4: Suboptimality Gap under Varying $\hat{\gamma}$ with $|\mathcal{D}|=30000$.
Figure 5: Suboptimality Gap under Varying $\hat{\gamma}$ with $|\mathcal{D}|=35000$.
Figure 6: Suboptimality Gap under Varying $\hat{\gamma}$ with $|\mathcal{D}|=40000$.
@@
C.3 The Dimension of Preference Vector
We first select $1k$ users and $1k$ items with the highest interaction frequencies, construct the binary feedback matrix based on users’ ratings. Then we use SVD to decompose the extracted binary feedback matrix $R_{u\times m}=\theta SX^{T}$, where $\theta=(\theta_{i}),i\in[u]$ and $X=(x_{j}),j\in[m]$. Note that the user $\bm{\theta}_{u}$ distribution in this setting does not fully match the assumptions from the simulation, as the user $\bm{\theta}_{u}$ values are scattered and exhibit clustering behavior.
We apply the Elbow Method to select the optimal dimension $d$. The objective is to identify the dimension that captures the most significant variance while ensuring computational efficiency. Figures 9 and 9 show the normalized variance explained by the first $d$ singular values for both datasets. In both cases, the "elbow" occurs at about $d=20$, where the marginal gain in explained variance diminishes. Thus, $d=20$ is selected as the optimal dimension for both datasets, achieving a balance between model complexity and variance preservation.
Figure 8: SVD Elbow Method for Yelp Dataset.
Figure 9: SVD Elbow Method for MovieLens Dataset.
C.4 User Probability Distribution (Detailed Analysis)
To determine the dataset size for each user, we consider two sampling strategies. When determining the scale of each $\mathcal{D}_{i}$ in offline data sampling, we mainly consider two distributions:
  • Equal Distribution: The probability of sampling data is uniform across all users, which gives
    $E(|\mathcal{D}_{i}|)=\frac{|\mathcal{D}|}{U}.$
  • Semi-Random Distribution: While the probability within each cluster is uniform, the probabilities of selecting different clusters vary. Each class has its own probability while users within the same class share the same probability. This can be expressed as
    $E(|\mathcal{D}_{i}|)=\frac{|\mathcal{D}|}{U}\cdot p_{J(i)},\quad p=[p_{1},p_{2},\ldots,p_{J}],\quad\sum_{i=1}^{J}p_{i}=1,\quad p_{i}\geq 0.$
@@ C.4 User Probability Distribution (Detailed Analysis)
DBSCAN_Improve 0.146196 0.069701 0.033431 0.018931 0.011499
XMeans_improve 0.187885 0.138490 0.085507 0.098608 0.142219
CLUB 0.133169 0.037698 0.021715 0.013878 0.010067
Off-CLUB 0.022586 0.021187 0.012889 0.007229 0.004594
LinUCB 0.152432 0.151958 0.152372 0.151933 0.152491
LinUCB_IND 0.143812 0.066350 0.027777 0.013595 0.007608
SCLUB 0.176953 0.121705 0.041028 0.040673 0.041705
Off-C2LUB_Pessimistic 0.021467 0.018053 0.011366 0.007157 0.004992
Off-C2LUB_Optimistic 0.023665 0.020053 0.011782 0.007168 0.004709
@@
OffDBSCAN_Improve 0.149179 0.072216 0.033121 0.015387 0.010090
OffXMeans_improve 0.200525 0.189838 0.178002 0.167926 0.159602
CLUB 0.147572 0.060904 0.026974 0.013053 0.007262
Off-CLUB 0.087765 0.042926 0.019609 0.010277 0.005976
LinUCB 0.159126 0.159954 0.159370 0.159623 0.159583
LinUCB_IND 0.145191 0.066178 0.028085 0.013468 0.007575
SCLUB 0.174606 0.117944 0.039357 0.039316 0.039230
Off-C2LUB_Pessimistic 0.083684 0.040423 0.019339 0.010653 0.006760
Off-C2LUB_Optimistic 0.089122 0.041868 0.019698 0.010243 0.005970
C.5 Offline Data Distribution (Detailed Analysis)
@@
The offline dataset is influenced by the user’s strategy during the sampling phase. For each data point in the dataset $\mathcal{D}_{u}$, a subset of 20 items ($\mathcal{A}_{i}$) is generated for sampling. We consider two selection strategies:
@@
  • Random Selection: The user randomly selects one item from $\mathcal{A}_{i}$, denoted as $a_{u}^{i}\sim\mathrm{Uniform}(\mathcal{A}_{i})$.
@@
  • LinUCB Selection: During the offline phase, each user independently applies the LinUCB method Abbasi-Yadkori et al. [2011] to select one item from $\mathcal{A}_{i}$, denoted as $a_{u}^{i}=\mathop{\mathrm{argmax}}\limits_{a\in\mathcal{A}_{i}}\langle\theta_{u},x_{a}\rangle+\alpha\sqrt{x_{a}^{\top}M_{u}^{-1}x_{a}}$.
When data is collected randomly, it implies that users are selecting items without exhibiting explicit preferences. Under such circumstances, the collected data spans all feature dimensions uniformly, ensuring broad coverage but lacking targeted exploration.
@@
By contrast, a more informed data collection strategy, such as LinUCB, allows the system to identify and concentrate on more informative dimensions. This offline process still relies on item-reward pairs from various arms, enabling the feature space to be explored in a more focused manner. Consequently, this approach yields tighter confidence intervals and enhances the estimation accuracy of the algorithm.
@@
In summary, while random selection ensures broad coverage of the feature space, LinUCB provides a more targeted exploration, leading to improved algorithmic performance. From a practical standpoint, if an existing recommendation system has been running under LinUCB, it can directly transition to the Off-C2LUB algorithm while leveraging the previously collected data. On the other hand, if newly designed algorithms demand uniform data distributions, it may be necessary to re-run the system in order to collect data adhering to these uniform assumptions.
C.6 Performance Analysis
Figure 7 illustrates the evolution of the average suboptimality gap under different real-world dataset scenarios (Yelp and MovieLens) over varying dataset sizes ($|\mathcal{D}|$). For Off-C2LUB, $\hat{\gamma}$ is optimized in an equal distribution scenario and applied to semi-random distribution. The results indicate that, despite no changes in the distribution of users’ $\bm{\theta}_{u}$, The optimized $\hat{\gamma}$ demonstrates environment sensitivity, necessitating adaptive strategies for robust performance. This underscores the importance of the optimistic-pessimistic strategy, which dynamically adapts to varying environments to achieve better performance. In real-world datasets, the distribution probability of users is their inherent attribute. In this case, the optimal gamma obtained through search can still achieve good performance under different sampling strategies. Notably, in the severely insufficient data regime ($|\mathcal{D}|\in[20k,100k]$), Off-C2LUB_Optimistic(Pessimistic) achieves significant improvements over baseline methods.
Table C.4 and Table C.4 present experimental suboptimality gap demonstrating that our proposed methods achieve excellent convergence, even in scenarios with relatively abundant data. Compared to traditional approaches such as CLUB and SCLUB, our methods exhibit faster convergence and maintain a significantly smaller suboptimality gap, highlighting their efficiency and robustness. This reduction in the suboptimality gap underscores the superior performance of our methods in achieving near-optimal solutions under varying data conditions.
Appendix D Further Discussions on Off-CLUB
D.1 Detailed Algorithm: Off-CLUB
The pseudo-code of Off-C2LUB is as follows:

### v2

@@ Appendix D Technical Lemmas
Lemma D.2.
Let $\bm{a}_{s}$ for $s\in[n]$ be samples generated under ˜3.1 and $M_{n}=\sum_{s=1}^{n}\bm{a}_{s}\bm{a}_{s}^{\top}$. Then event
$\lambda_{\min}(M_{n})\geq\frac{1}{2}\tilde{\lambda}_{a}n,\ \forall n\geq\frac{16}{\tilde{\lambda}_{a}^{2}}\log\left(\frac{8d}{\tilde{\lambda}_{a}^{2}\delta}\right)$
holds with probability at least $1-\delta$.
Lemma D.3.
Let $\bm{a}_{s}$ for $s\in[n]$ be samples generated under ˜A.1 and $M_{n}=\sum_{s=1}^{n}\bm{a}_{s}\bm{a}_{s}^{\top}$. Then event
@@
$\lambda_{\min}(M_{n})\geq 2\lambda_{a}n>\frac{1}{2}\lambda_{a}n,\ \forall n\geq\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8d}{\lambda_{a}^{2}\delta}\right)$
@@
When $n\geq\frac{512d}{\gamma^{2}\tilde{\lambda}_{a}}\log(\frac{2U}{\delta})$ and $\lambda\leq d\log(1+\frac{n}{\lambda d})+2\log\left(\frac{2U}{\delta}\right)$,
$\frac{\sqrt{d\log(1+\frac{n}{\lambda d})+2\log(\frac{2U}{\delta})}+\sqrt{\lambda}}{\sqrt{\tilde{\lambda}_{a}\lambda+\tilde{\lambda}_{a}n/2}}\leq\frac{\gamma}{4}.$
@@
The proofs of the above lemmas can be found in Li and Zhang [2018] and Wang et al. [2025]. Specifically, Lemma D.1 corresponds to Lemma 1, Lemma D.4 to Lemma 9, and Lemma D.5 to Lemma 10 in Li and Zhang [2018]. Furthermore, Lemma D.2 comes from combining ˜3.1, Lemma J.1 in Wang et al. [2023a] and Lemma 7 in Li and Zhang [2018], while Lemma D.3 is from Lemma 2 in Li et al. [2025], Lemma 7 in Li and Zhang [2018] and Lemma B.2 in Wang et al. [2025].
Appendix E Additional Experiments
E.1 Basic Setup and Dataset Description
Dataset. We fix the number of users at $U=1\text{k}$ and the preference vector dimension at $d=20$. The dataset size $|\mathcal{D}|$ ranges from $5\text{k}$ to $100\text{k}$ in the insufficient-data setting and from $0.2\text{M}$ to $1\text{M}$ in the sufficient-data setting. For each experiment, the first half of $|\mathcal{D}|$ is used for offline training and the second half for evaluation. All experiments are repeated 10 times with random seeds, with parameters fixed as $\lambda=0.5$, $\alpha=0.1$, and $\delta=0.01$. For ARMUL, we set the learning rate $\eta=0.01$, the number of iterations $T_{\text{iteration}}=10$, and performed an $n_{\text{fold}}=2$ uniform split of the dataset.
@@ E.1 Basic Setup and Dataset Description
(1) Synthetic Dataset: User preference vectors $\bm{\theta}_{u}$ and item feature vectors $\bm{a}_{u}^{i}$ lie in a $d=20$-dimensional space, with entries drawn from a standard Gaussian distribution and normalized to unit norm. Users are uniformly partitioned into $J=10$ clusters. For each user, the candidate action set $\mathcal{A}_{u}$ contains 20 items, whose feature vectors are sampled independently from a Gaussian distribution and normalized. Rewards are simulated as
$r_{u}^{i}=\langle\bm{\theta}_{u},\bm{a}_{u}^{i}\rangle+\epsilon_{u}^{i},$
where $\epsilon_{u}^{i}$ is zero-mean sub-Gaussian noise with variance parameter $\sigma=0.05$.
(2) Real-World Datasets: We select the top 1,000 users and items by interaction frequency and construct the user–item feedback matrix. User preference vectors are extracted via Singular Value Decomposition (SVD) following Li et al. [2019], i.e.,
$R_{U\times M}=\Theta SX^{\top},\quad\Theta=(\bm{\theta}_{u})_{u\in\mathcal{U}},\quad X=(x_{j})_{j\in[M]}.$
In this setting, user preferences are scattered and exhibit clustering behavior, which differs from the synthetic assumption. Each user’s candidate action set $\mathcal{A}_{u}$ again contains 20 items, generated by sampling sub-Gaussian feature vectors and normalizing them to unit norm.
User Probability Distribution. To allocate the dataset size across users, we consider two sampling strategies:
  • Equal Distribution: Each user is sampled with equal probability, i.e.,
    $\mathbb{E}[|\mathcal{D}_{u}|]=\frac{|\mathcal{D}|}{U}.$
  • Semi-Random Distribution: Sampling is uniform within each cluster, but probabilities differ across clusters. Let $j_{u}$ denote the cluster index of user $u$, and $\mathcal{V}(j_{u})$ the set of users in cluster $j_{u}$. Then
    $\mathbb{E}[|\mathcal{D}_{u}|]=\frac{|\mathcal{D}|}{|\mathcal{V}(j_{u})|}\cdot p_{j_{u}},\quad p=[p_{1},\ldots,p_{J}],\quad\sum_{j=1}^{J}p_{j}=1,\quad p_{j}\geq 0.$
E.2 Benchmark Methods and Their Descriptions
To ensure fairness, we simulate online algorithms (e.g., CLUB) in the offline setting by treating the dataset as a one-shot interaction stream. Each algorithm runs its original policy to make predictions without further exploration. Clustering is done once before the evaluation stage, avoiding dynamic updates and preserving the offline nature. It is important to note that, without modification, the original method may lead to meaningless updates during offline training. Specifically, in the case of CLUB, removing edges under insufficient information can cause irreversible loss of structural information.
Baseline Algorithms. To ensure comprehensive and fair comparisons, we select the following baseline algorithms from existing studies. We adapt several online learning algorithms to the offline setting, modifying them as necessary to work effectively with offline datasets. Additionally, we propose enhanced versions of DBSCAN [Schubert et al., 2017], specifically tailored to our experimental environment. For these multi-armed bandit (MAB) methods, we first use offline training data to update model parameters, such as the preference vector $\bm{\theta}$. We then refine the graph or set structure based on the updated parameters.
For offline clustering algorithms. We incorporate DBSCAN_Improve and XMeans_Improve to replace the clustering mechanism in CLUB. Furthermore, we uniformly augment all compared methods by introducing an upper confidence bound (UCB) term to ensure fair comparison. The evaluated baseline methods include:
@@ E.2 Benchmark Methods and Their Descriptions
  • LinUCB_Ind [Abbasi-Yadkori et al., 2011]: An independent variant of LinUCB that operates without leveraging any clustering information. This method applies the LinUCB approach independently to each user.
@@
  • CLUB [Gentile et al., 2014]: A cluster-based multi-armed bandit method that groups users from a complete graph according to their characteristics to improve model selection.
  • DBSCAN_Improve [Schubert et al., 2017]: A variant of CLUB that replaces its original clustering mechanism with the DBSCAN algorithm.
  • XMeans_Improve [Pelleg et al., 2000]: A variant of CLUB that replaces its original clustering mechanism with the X-Means algorithm.
  • SCLUB [Li et al., 2019]: A method that employs a hierarchical clustering structure to model users and enhances learning efficiency by optimizing cluster centers.
  • Off-CLUB (Algorithm˜2): A cluster-based multi-armed bandit method that groups users from a complete graph according to their characteristics to improve model selection. It utilizes a Lower Confidence Bound (LCB) exploration strategy and performs aggregation through neighbor information.
  • Off-C2LUB (Algorithm˜1): A cluster-based multi-armed bandit method that groups users from a null graph based on their characteristics to enhance model selection. It employs an LCB exploration term and aggregates feedback via neighboring nodes. The $\hat{\gamma}$ is selected to be approximate to the optimal value.
  • Off-C2LUB_Underestimation: A variant of Off-C2LUB that uses an underestimation-based estimation policy to compute $\hat{\gamma}$.
  • Off-C2LUB_Overestimation: A variant of Off-C2LUB that employs an overestimation-based estimation policy to determine $\hat{\gamma}$.
  • ARMUL [Duan and Wang, 2023]: A framework named Adaptive and Robust MUlti-task Learning (ARMUL) addresses the multi-task learning problem. Use a regularization term to balance the similarity and differences between tasks.
Figure 4: Comparison of clustering strategies under $|\mathcal{D}|=30k$ across three datasets.
Figure 5: Comparison of clustering strategies under $|\mathcal{D}|=35k$ across three datasets.
Figure 6: Comparison of clustering strategies under $|\mathcal{D}|=40k$ across three datasets.
@@
E.3 Offline Data Distribution (Detailed Analysis)
@@
DBSCAN_Improve 0.146616 0.069673 0.033521 0.018761 0.011276
XMeans_Improve 0.178563 0.138018 0.103209 0.089212 0.086906
CLUB 0.144821 0.066241 0.028270 0.013642 0.007569
Off-CLUB 0.007643 0.002359 0.001360 0.000968 0.000751
LinUCB_IND 0.144928 0.066318 0.028251 0.013649 0.007584
SCLUB 0.097121 0.045636 0.020032 0.010014 0.005680
ARMUL 0.137016 0.112671 0.095382 0.085142 0.070668
Off-C2LUB 0.010966 0.009098 0.008366 0.003407 0.000778
Off-C2LUB_Underestimation 0.007720 0.002684 0.001758 0.001340 0.001126
Off-C2LUB_Overestimation 0.007579 0.002353 0.001345 0.000950 0.000751
@@
DBSCAN_Improve 0.150422 0.071688 0.031250 0.016067 0.009879
XMeans_Improve 0.202409 0.190408 0.178008 0.168371 0.160795
CLUB 0.144998 0.066608 0.028231 0.013636 0.007575
Off-CLUB 0.010777 0.002992 0.001580 0.001077 0.000818
LinUCB_IND 0.145213 0.066277 0.028228 0.013666 0.007573
SCLUB 0.140205 0.064720 0.027545 0.013374 0.007443
ARMUL 0.119995 0.089696 0.091467 0.078193 0.057992
Off-C2LUB 0.016508 0.014446 0.013647 0.013440 0.013294
Off-C2LUB_Underestimation 0.011465 0.004185 0.002810 0.002353 0.002099
Off-C2LUB_Overestimation 0.010776 0.002980 0.001589 0.001081 0.000821
@@
The offline dataset is influenced by the user’s strategy during the sampling phase. For each data point in the dataset $\mathcal{D}_{u}$, a subset of 20 items ($\mathcal{A}_{u}$) is generated for sampling. We consider two selection strategies:
@@
  • Random Selection: The user randomly selects one item from $\mathcal{A}_{u}$, denoted as $\bm{a}_{u}^{i}\sim\mathrm{Uniform}(\mathcal{A}_{u})$. Figure 2 shows the experimental results.
@@
  • LinUCB Selection: During the offline phase, each user independently applies the LinUCB method [Abbasi-Yadkori et al., 2011] to select one item from $\mathcal{A}_{u}$, denoted as $\bm{a}_{u}^{i}=\mathop{\mathrm{argmax}}\limits_{\bm{a}\in\mathcal{A}_{u}}\langle\bm{\theta}_{u},\bm{a}\rangle+\alpha\sqrt{\bm{a}^{\top}M_{u}^{-1}\bm{a}}$. Figure 7 shows the experimental results.
@@
When data is collected randomly, it suggests that users are choosing items without demonstrating explicit preferences. In this case, the collected data covers all feature dimensions in a relatively uniform manner, ensuring broad coverage but lacking targeted exploration. In contrast, a more informed collection strategy, such as LinUCB, enables the system to identify and focus on more informative dimensions. Although this process still relies on item–reward pairs from different arms, it guides the exploration of the feature space in a more deliberate way. As a result, the algorithm achieves tighter confidence intervals and improved estimation accuracy.
@@
In summary, while random selection guarantees comprehensive coverage of the feature space, LinUCB offers more targeted exploration, ultimately enhancing algorithmic performance. From a practical perspective, if a recommendation system has already been operating under LinUCB, it can seamlessly transition to the Off-C2LUB algorithm while reusing previously collected data. Conversely, if a newly designed algorithm requires uniformly distributed data, it may be necessary to re-run the system to gather data consistent with such assumptions.
E.4 Performance Analysis
Performance with Insufficient Dataset. To illustrate the performance under insufficient data with LinUCB selection, we report the average suboptimality gap across user distributions and dataset sizes (Figure 7). All results are averaged over 10 random runs, and the error bars are computed using standard errors. For Off-C2LUB, $\hat{\gamma}$ is optimized under the equal distribution scenario and applied to the semi-random distribution. With $|\mathcal{D}|\in[20k,100k]$, both Off-C2LUB variants significantly reduce the suboptimality gap. On the synthetic dataset, Off-C2LUB_Overestimation improves over Off-CLUB by 45.4% and over other baselines by at least 76.3%, while Off-C2LUB_Underestimation achieves 33.7% and 71.3%, respectively. On Yelp, the corresponding improvements are 27.2% and 76.5% for Overestimation, and 21.4% and 74.6% for Underestimation. On MovieLens, they are 17.2% and 75.3% for Overestimation, and 14.4% and 63.4% for Underestimation. Overall, Off-C2LUB achieves the best performance, while Off-CLUB, though less effective, still outperforms classical clustering baselines.
Performance with Sufficient Data. Table E.3 and Table E.3 further report the convergence performance in the real dataset under sufficient data. The bold values mark the optimal suboptimal gaps (rounded to six decimals) for each dataset size. Our proposed Off-CLUB and Off-C2LUB consistently outperform all other baselines. Compared to classical methods such as CLUB and SCLUB, our approaches converge faster and consistently maintain a much smaller suboptimality gap, underscoring their efficiency and robustness.
Improvement with Estimated $\hat{\gamma}$. We evaluate different strategies under limited data conditions using the synthetic, Yelp, and MovieLens datasets with $|\mathcal{D}|=30k$ (Figure 4), $|\mathcal{D}|=35k$ (Figure 5), and $|\mathcal{D}|=40k$ (Figure 6). Solid lines denote methods using $\hat{\gamma}$, and dashed lines those without. Consistent with the main text, our algorithm achieves near-optimal performance compared to the Off-C2LUB lower bound. This result holds across different dataset scales and domains, regardless of the $\hat{\gamma}$ estimation method, demonstrating that both estimation strategies maintain stable and competitive performance.
Appendix F Further Discussions on Off-CLUB
F.1 Detailed Algorithm: Off-CLUB
The pseudo-code of Off-C2LUB is shown in Algorithm˜2. The algorithm takes similar inputs to Algorithm 1, except that it does not require $\hat{\gamma}$. It starts with a complete graph $\mathcal{G}$ and iterates through all edges $(u_{1},u_{2})\in\mathcal{E}$ during the Cluster Phase. If the condition $\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}>\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right)$ holds, the algorithm concludes that users $u_{1}$ and $u_{2}$ do not share the same preference vector and removes the edge between them (Line 5). After processing all edges, the resulting graph $\tilde{\mathcal{G}}$ separates users with different preference vectors into distinct clusters, provided that each user has sufficient samples to ensure a low confidence interval. However, if a user $u$ has fewer samples in $\mathcal{D}_{u}$, the resulting higher confidence interval $\text{CI}_{u}$ reduces the likelihood of satisfying condition (6), potentially leading to the inclusion of heterogeneous users in the same cluster. The conditions necessary to ensure the algorithm’s correctness are discussed in detail in Section 5.2.